# Cooperation as Black Box: Conceptual Fluctuation and Diagnostic Tools for Misalignment in MAS

### v1

@@ 4 Discussion
MAS capabilities are widely acknowledged, as their architectures and applications offer unique advantages, which are particularly appealing for dynamic and distributed systems, enabling the efficient management of complex tasks and environments [48]. For example, Liu et al.[49] emphasize the promise of MAS in optimizing the scheduling of aggregated resources and solving problems dynamically, demonstrating the transformative potential of these systems across domains.
Indeed, these technologies are already in active deployment, as noted by Müller et al. [50]. Thus, Xue et al. [51] describe a MAS-based control strategy for managing individual entities and bottom-layer control systems, highlighting MAS’s potential to revolutionize industries through improved coordination, scalability, and adaptability. However, as deployment scenarios become more complex, the distinction between coordination and cooperation becomes operationally indistinct under real-world conditions. Without standardized frameworks, MAS models risk misclassifying agent behaviors, leading to inefficiencies and unintended emergent properties.
The Rabbit-Duck illusion exemplifies how the same visual input can yield different interpretations. Similarly, agent behaviors in MAS can be perceived as either cooperative or only coordinative, depending on theoretical and methodological perspectives. Therefore, the ambiguity between coordination and cooperation complicates MAS evaluation, necessitating clearer guidelines and standardized definitions. Recent work by Kierans et al. [52] highlights the need to frame alignment challenges from a sociotechnical perspective, incorporating multiple human and AI agents. While our focus remains on conceptual instability within MAS, their call reinforces that misalignment is layered — emerging not only from technical failures, but from interpretive ambiguity embedded in system design and evaluation.
We introduced the Misalignment Mosaic, a diagnostic framework that articulates how overlooked semantic and interpretative distinctions (especially between cooperation and coordination) contribute to misalignment in MAS. We present below the implications of the Misalignment Mosaic:
@@
    Recognizing misalignment as a multi-layered issue makes it clear that no single intervention can fully resolve systemic failures. A mosaic perspective compels experts to assess the entire ecosystem of MAS (spanning from conceptual definitions, benchmarking practices, and real-world interpretability) rather than treating misalignment as a distant/abstract problem.
@@
    Concept to Code Decay and Early-Stage Alignment Checks: Conceptual models may lose key properties during implementation and deployment, resulting in different representativeness. This distorts intended functionality and introduces ethical concerns, especially in cases where AI systems interact with humans. Therefore, by viewing misalignment as a network of interlinked issues, designers are encouraged to integrate conceptual clarity into early development stages. This means explicitly defining coordination and cooperation before they manifest in code, ensuring that user-facing interfaces reflect intended system behaviors rather than assumed ones.
@@
    Refinement of terminology and benchmarks: without structured documentation and classification, the MAS community faces significant barriers to aligning with both technical and moral goals. The mosaic framework highlights the need for standardized definitions and clear evaluation frameworks that distinguish coordination from cooperation. Without precise benchmarks, both experimental results and real-world deployments risk misinterpreting system alignment and, therefore, leading to inefficiencies, failure cases, or unverified claims about agent behavior.
@@
    Magnification: identified challenges are now scaling to agentic workflows, where coordination and cooperation ambiguities become even more important. Large-scale industrial deployments demand trust, reliability, and interpretability.
Call for action. We propose a multidisciplinary effort led by the MAS community. We urge the community to partner with other disciplines and undertake a systematic investigation of terminology across the literature and: 1) Develop a standardized MAS taxonomy for coordination, cooperation, and variants. 2) Create dynamic platforms and benchmarks for evaluating those under varied conditions. These platforms would provide the necessary scaffolding to assess and improve coordination and cooperation mechanisms. Finally, 3) Integrate sociotechnical perspectives: expanding alignment research to humanity-centered design considerations. This reinforces the idea that cooperation and coordination are not just technical constructs but integral to both moral and functional systems [42], and key to other disciplines as well.
5 Conclusion: Misalignment Begins with Meaning
Starting from the identification of three interconnected drivers of misalignment in MAS (coordination–cooperation ambiguity, conceptual fluctuation, and semantic instability), we introduced the Misalignment Mosaic as a framework for diagnosing and tracing interpretive misalignment in MAS. The coordination/cooperation ambiguity is not a mere terminological detail, but a foundational fault line that permeates the design, evaluation, and governance of MAS and AI systems more broadly.
This work does not propose an empirical resolution to this ambiguity, but rather calls for a structural shift in how the MAS community frames, audits, and addresses meaning-level misalignment. Conceptual clarity must precede empirical validation; as long as foundational MAS concepts remain unstable, no amount of technical refinement will ensure alignment at scale.
While our discussion draws from existing empirical studies in MAS and MARL, we emphasize that the next step is not immediate experimentation, but a coordinated, interdisciplinary effort to establish precise terminology, benchmarks, and conceptual scaffolding. Clarifying meaning is a prerequisite for building aligned, trustworthy AI. Finally, if we seek to build AI systems that cooperate purposefully with humans and across cultural boundaries, it is vital to establish shared conceptual scaffolds that define cooperation across disciplinary, institutional, and deployment contexts.

### v2

@@ 4 Discussion
MAS capabilities are widely acknowledged, as their architectures and applications offer unique advantages, which are particularly appealing for dynamic and distributed systems, enabling the efficient management of complex tasks and environments [kalachev2018intelligent]. For example, Liu et al.[liu2019multi] emphasize the promise of MAS in optimizing the scheduling of aggregated resources and solving problems dynamically, demonstrating the transformative potential of these systems across domains.
Indeed, these technologies are already in active deployment, as noted by Müller et al. [muller2014application]. Thus, Xue et al. [xue2023rapid] describe a MAS-based control strategy for managing individual entities and bottom-layer control systems, highlighting MAS’s potential to revolutionize industries through improved coordination, scalability, and adaptability. However, as deployment scenarios become more complex, the distinction between coordination and cooperation becomes operationally indistinct under real-world conditions. Without standardized frameworks, MAS models risk misclassifying agent behaviors, leading to inefficiencies and unintended emergent properties.
The Rabbit-Duck illusion exemplifies how the same visual input can yield different interpretations. Similarly, agent behaviors in MAS can be perceived as either cooperative or only coordinative, depending on theoretical and methodological perspectives. Therefore, the ambiguity between coordination and cooperation complicates MAS evaluation, necessitating clearer guidelines and standardized definitions. Recent work by Kierans et al. [kierans2024] highlights the need to frame alignment challenges from a sociotechnical perspective, incorporating multiple human and AI agents. While our focus remains on conceptual instability within MAS, their call reinforces that misalignment is layered, emerging not only from technical failures but from interpretive ambiguity embedded in system design and evaluation.
We introduced the Misalignment Mosaic, a diagnostic framework that articulates how overlooked semantic and interpretative distinctions (especially between cooperation and coordination) contribute to misalignment in MAS. We list below some implications of the Misalignment Mosaic:
@@
    Recognizing misalignment as a multi-layered issue makes it clear that no single intervention can fully resolve systemic failures. A mosaic perspective compels experts to assess the entire ecosystem of MAS (spanning from meaning and conceptual definitions, benchmarking practices, and real-world interpretability) rather than treating misalignment as a distant problem.
@@
    Concept-to-Code Decay and Early-Stage Alignment Checks: conceptual models may lose key properties during implementation and deployment, resulting in reduced representativeness. This distorts the intended functionality and raises ethical concerns, especially when algorithmic systems interact with humans. Therefore, by viewing misalignment as a network of interlinked issues, designers are encouraged to integrate conceptual clarity into the early stages of development. This means explicitly defining coordination and cooperation before they manifest in code, ensuring that user-facing interfaces reflect intended system behaviors rather than assumed ones.
@@
    Refinement of terminology and benchmarks: without structured documentation and classification, the MAS community faces significant barriers to aligning with both technical and moral goals. The Misalignment Mosaic underscores the necessity for standardized definitions and clear evaluation frameworks that differentiate coordination from cooperation. Otherwise, both experimental results and real-world deployments risk misinterpreting system alignment and, therefore, leading to inefficiencies, failure cases, or unverified claims about agent behavior.
@@
    Urgency: identified challenges are now scaling to agentic workflows, where coordination and cooperation ambiguities become even more dangerous. Large-scale industrial deployments demand trust, reliability, and interpretability.
Call for action. We urge the community to partner with other disciplines and undertake a systematic investigation of terminology across the literature and: 1) Develop a standardized MAS taxonomy for coordination, cooperation, and variants. 2) Create dynamic platforms and benchmarks for evaluating those under varied conditions. These platforms would provide the necessary scaffolding to assess and improve coordination and cooperation mechanisms. Finally, 3) Integrate sociotechnical perspectives: expanding alignment research to multidisciplinary and multicultural considerations. This reinforces the idea that cooperation and coordination are not just technical constructs but integral to both moral and functional systems [curry2019good], and vital to other disciplines as well.
Limitations and Future Directions. This work adopts a conceptual and diagnostic approach to misalignment in MAS, focusing on the meaning-level and interpretive conditions that precede technical failures. While we present the Misalignment Mosaic, we must acknowledge its limitations: our analysis centers on the ambiguity between cooperation and coordination. While we argue that this ambiguity is foundational to many misalignment risks, other overloaded terms in MAS (such as autonomy, alignment, or trust) require parallel diagnostic attention. Future efforts should explore how semantic instability propagates across AI domains, particularly as MAS-like architectures underpin increasingly complex agentic systems.
Despite these limitations, we believe this work offers a necessary foundation for auditing meaning as a structural dimension of alignment, a dimension often overlooked in technical models of MAS design and governance.
5 Conclusion
We introduced the Misalignment Mosaic as a framework for diagnosing and tracing meaning-level misalignment in MAS. This work does not propose an empirical resolution to this ambiguity, but rather calls for a structural shift in how the MAS community frames, audits, and addresses meaning-level alignment. Conceptual clarity must precede empirical validation; as long as foundational MAS concepts remain unstable, no amount of technical refinement will ensure alignment at scale.
While our discussion draws from existing empirical studies in MAS and MARL, we emphasize that the next step is not immediate experimentation, but a coordinated, multidisciplinary effort to establish clearer terminology, benchmarks, and conceptual scaffolding. Clarifying meaning is a prerequisite for building aligned, trustworthy algorithmic systems.
We invite the community to reflect: is there a connection between misalignment and moral overreadings with treating cooperation as a black box? If we seek to build AI systems that cooperate purposefully with humans and across cultural boundaries, it is vital to establish shared conceptual scaffolds that define cooperation across disciplinary, institutional, and deployment contexts.